OpenStudy (anonymous):

Phillip bought 12 used CDs and DVDs. CDs cost $2 each and DVDs cost $3 each. He spent $31 not including tax. How many DVDs did Phillip buy?

OpenStudy (cgreenwade2000):

I love explaining these. Are you ready to learn sir/madam?

OpenStudy (anonymous):

Yes

OpenStudy (cgreenwade2000):

Alright. There are 2 unknowns here correct? Tell me what they are.

OpenStudy (anonymous):

how many cds he bought and how many dvds he bought I'm guessing

OpenStudy (cgreenwade2000):

Yes. For now, we'll call the number of CDs by the variable, C. and the number of DVDs will be D.

OpenStudy (anonymous):

Alright sounds good

OpenStudy (cgreenwade2000):

Using those variables, express the total number of CDs and DVDs he bought.

OpenStudy (anonymous):

12

OpenStudy (cgreenwade2000):

Use the variables in an equation.

OpenStudy (anonymous):

C + D = 31

OpenStudy (cgreenwade2000):

Incorrect. What we want is (the number of CDs) + (the number of DVDs) = 12

OpenStudy (anonymous):

2C + 3D = 12

OpenStudy (cgreenwade2000):

You are getting ahead of yourself.

OpenStudy (cgreenwade2000):

What is the variable for (number of CDs)?

OpenStudy (anonymous):

C?

OpenStudy (cgreenwade2000):

Yes. We already named it. So if we know (the number of CDs) + (the number of DVDs) =12 Write an equation using variables.

OpenStudy (anonymous):

C + D = 12?

OpenStudy (cgreenwade2000):

Correct. That's the first part of the story problem. Now for part 2! How much does a CD cost? How much does a DVD cost?

OpenStudy (anonymous):

2$ and 3$

OpenStudy (cgreenwade2000):

So if he bought 3 CDs what would the price be?

OpenStudy (anonymous):

6$

OpenStudy (cgreenwade2000):

Good. And we get that by (the price of a CD)x(the number of CD) or 2C

OpenStudy (cgreenwade2000):

So let's write an equation to find the total price of everything altogether. We know it's $31

OpenStudy (anonymous):

Right so 2C + 3D = 31?

OpenStudy (cgreenwade2000):

Yes! Now we have 2 equations C+D=12 and 2C + 3D =31

OpenStudy (cgreenwade2000):

Now for the fun stuff. If C + D =12 , then what does D equal?

OpenStudy (anonymous):

3?

OpenStudy (cgreenwade2000):

Don't guess.

OpenStudy (cgreenwade2000):

If C + D = 12 then D = 12 - C

OpenStudy (cgreenwade2000):

TO get that, we would just subtract C on both sides of the equal sign.

OpenStudy (anonymous):

ok got it

OpenStudy (cgreenwade2000):

Alright so now we know what D equals. We want to substitute it into the second equation. 2C + 3D = 31 2C + 3(12-C) =31

OpenStudy (anonymous):

2C-36+3C=31 2C+3C=31+36 5C=67

OpenStudy (cgreenwade2000):

You have to do that one again. I accidentally wrote it incorrectly.

OpenStudy (cgreenwade2000):

It is +3D so it should be 3(12-C)

OpenStudy (anonymous):

so wait the 2D is in front so it should be 2D + 3(12-7) = 31?

OpenStudy (cgreenwade2000):

Where are you getting 7 from? You should be working on solving this 2C + 3(12-C) =31

OpenStudy (anonymous):

whoops sorry i meant c

OpenStudy (cgreenwade2000):

Go ahead and solve that equation for C

OpenStudy (anonymous):

2C + 3(12-C) =31 2C + 36 - 3C = 31 2C - 3C = 31-36 -1C=-5

OpenStudy (cgreenwade2000):

Finish solving. Right now you have -C=-5 Solve so it says C=????

OpenStudy (anonymous):

5

OpenStudy (cgreenwade2000):

:D So C=5

OpenStudy (cgreenwade2000):

So the number of CDs is what?

OpenStudy (anonymous):

5

OpenStudy (cgreenwade2000):

Alright. Knowing what C is, can you solve for D?

OpenStudy (anonymous):

C+D=12 5+D=12 D=12-5 D=7

OpenStudy (cgreenwade2000):

So the answer that the problem asks you for is??????

OpenStudy (anonymous):

How many DVDs did Phillip buy. which is 7

OpenStudy (cgreenwade2000):

Congrats! Feel free to check your answer by plugging in 5 for C and 7 for D in the second equation.

OpenStudy (anonymous):

Yay thank you very much for your help :)